f(x)=x^2+2x+1 minimum or maximum value?
Found 2 solutions by Alan3354, fractalier:
Answer by Alan3354(69443)

The coefficient of the x^2 is + --> minimum.

Answer by fractalier(6550)

A couple of ways to tell...
Since it curves concave up, it has a minimum...x = -b/2a = -2/2 = -1
Second derivative test...first derivative is 2x + 2...second derivative is 2 and always positive, hence it has a minimum...